January 19, 2011 Tracking CDC/FDC

From GlueXWiki
Jump to: navigation, search

Meeting Time and Place

Wednesday January 19, 2011 at 11:00am At Jefferson Lab, the meeting will be held in F326


To connect from the outside, please use ESNET
1.) ESNET: 8542553
2.) Phone:
+1-800-377-8846 : US
+1-888-276-7715 : Canada
+1-302-709-8424 : International
then enter participant code: 77438230# (remember the "#")


  1. Announcements
  2. The Collaboration Meeting
  3. FDC Update logbook
  4. CDC Update [1]
  5. Tracking Issues


Participants: Curtis, Naomi, Beni, Simon, Lubomir, Dave and Mark

Collaboration Meeting

Curtis has scheduled one talk each for the FDC and CDC groups. The next section is Offline Reconstruction which will follow on nicely, there will probably be an overview by Mark and reports on Dmitri's calibration database and tracking status. Beni and Lubomir will let Curtis know in the next few days if they need to make another talk on the FDC.

FDC Update

  • Clean room: final walk-through with the vendor; will move in by the end of January.
  • We have all the parts and materials needed to start the chamber production. First article rigid-flex assemblies were inspected but for the mass production we need to decide about the tinning of the connectors (depending on the method of soldering). First article signal cables have one of the two connectors made with opposite polarity, other than that the quality is acceptable. They will be fixed by the vendor.
  • We are testing our new stringing tools exercising different procedures. We measured the tensions of an old full-scale prototype wire frame and compared them to the measurements done few years ago when the frame was made at IUCF (FDC logbook page 548). Apart from a systematic shift by ~7% there's very good agreement wire by wire. It means 1) our measurements are reliable, and 2) the frame didn't deteriorate during these years. We strung wires for a full chamber and measured their positions using laser sensor and automated stepper motor (page 549). Position resolution of ~10-15 microns was achieved. The main issue now is the flatness of the wire frame. The deformation of the frame results in gaps between the wires and soldering pads.

CDC Update

Curtis reported on the CDC construction status:

  • the remaining straws in row 10 will be put into place today and glued tomorrow. Photos of CDC We are completing about one layer each week. An undergrad student is working on cutting more straws to size and others are checking the parts - there are still a few 1000 pinholders to be checked and cleaned. Installing the straws takes 3 people, usually these are the 2 techs and Gary; Curtis and Naomi have been trained for use if needed.
  • We have not found any more bad resistance straws - these all came from one section of the box. Initially we found 2 straws where where the straw-donut glue-joint made no electrical connection, this has not been seen since then. The resistance of each straw and straw-donut is being checked.
  • Gary has made some modifications to Slava's test electrical hookup. Photos of hookup tests He found that it can be sealed with less torque than initially specified. He also found that the wires leaked gas (along the mesh) but can sealed with cyanoacrylic over the braid and centre conductor before the heatshrink is applied. The glue wicks up the cable. We plan to train students to prepare the wires - cut, strip, solder ball, seal, add the conductive rubber and heatshrink, and then make sure that the finished wires have low resistance.
  • Our particulate meter was returned from servicing and indicates that we still have a class 2000 cleanroom. The low humidity has been causing static electricity problems (eg giving people a small shock when they touch the table) so we have been trying to raise humidity and also grounded the cleanroom and the table to the building ground. After installing the straws in the CDC, the ends are sealed with a plastic cap to prevent dust from collecting inside them.

Naomi has been installing Xilinx software so that she can download software from Gerard into the new fADC to increase the pre-trigger buffer time. She is preparing to send some of our MFCs back to MKS for recalibration.

Tracking Issues

Dave Lawrence spoke about some issues in the tracking software...

  1. In the reconstruction software certain events would hang or segfault, but not reproducibly, on the IU 32-bit Linux farm. This might be associated with the fact that the code version combined with the machine/OS caused them to be using the now deprecated ALT1 least-squares fitter rather than the KalmanSIMD. Since the problem can be avoided by minor and seemingly unrelated changes in the code, it cannot be determined if the ALT1 fitter is actually responsible for or merely exposing the problem. Work continues to try and identify the source of this.
  2. Different mass hypotheses for the same track (eg fit pi+ to both p and pi+) have been found to give a large difference in reconstructed momentum, possibly due to a different hit selection. Simon is looking at activating the Kalman filter to reject certain hits and Dave is looking at changing the CDC hit selection, the sigma calculation looked wrong... this will be ongoing for the next few days.
  3. Historically -999 in code has been used to indicate not a valid number, some of these were propagating by division into a sensible number. Dave is now changing these to NAN. So... in future we will need to check if something is NAN before operating on it.

The software workshop had a visitor Dmitri Emeliyanov from the ATLAS tracking group at RAL. The ATLAS tracking process takes 1ms whereas the estimates for GlueX are around 20ms. Simon made some changes based on this. He found that the stepsize is too small and there are too many iterations/path so he doubled the stepsize and decreased iterations, now down to 10ms with no noticeable effect on resolution. Dave has added hi-res timers. Checking material maps for boundaries adds time. Other ATLAS characteristics are very similar, a 2T field track, lower measurement range of 100MeV, momentum resolution of a few %. Dave and Simon are going to look at probabilistic fitting, including both the left and right solutions simultaneously while fitting tracks.

Curtis said that Will found that some errors in the kinematic fitter were possibly too small, he will report on this at the collaboration meeting.

Next meeting

will be in 4 weeks, Feb 16.